z-4 over z-2 equals 1 over z-2
Answer by checkley77(12844)

(z-4)/(z-2)=1/(z-2) CANCEL OUT THE DENOMINATORS.
Z-4=1
Z=1+4
z=5 ans.
Proof:
(5-4)/(5-2)=1/(5-2)
1/3=1/3